OpenStudy (anonymous):

Can you form a right triangle with the three lengths given? Give a yes/no after showing work. a = 15 b = 25 c = 20 (Show Work)

jimthompson5910 (jim_thompson5910):

Check to see if x^2+y^2 = z^2 is true where 'z' is the longest length x^2 + y^2 = z^2 15^2 + 20^2 = 25^2 225 + 400 = 625 625 = 625 Since the last equation is true, this means that we can indeed form a right triangle with these given lengths.